the difference of two numbers is 9. twice the first number added to the second number is 24. find the two numbers.
Found 2 solutions by ewatrrr, Jk22:
Answer by ewatrrr(24785)


Hi
difference of two numbers is 9
Let x and (x+9) represent the first and second number respectively
Question states***
2x + (x+9) = 24
Solving for x
3x = 15
x = 5 Second number is 14 (5+9)
CHECKING our Answer***
10 + 14 = 24

Answer by Jk22(389)


let a,b be the 2 numbers, then

a-b=9
2a+b=24

from the 1st equation we get a=9+b, if we substitute in the 2nd :
18+2b+b=24 hence 3b=6, and so b=2, giving a=11

we verify : 11-2=9 and 22+2=24